Fermi National Laboratory


Accelerator Update for August 24, 2005 - August 26, 2005

Wednesday, August 24
The day shift began with Operations monitoring store 4347 and the Recycler stash, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using colliding beam, and with MiniBooNE, NuMI, and the Meson MCenter experiments taking beam.

Two Main Injector (MI) experts began a slipstacking study at 9:44 AM. They ended the study at 10:53 AM due to beam intensity variations.

Tevatron (TeV) experts corrected the orbits at sectors E11 and C49 and improved the beamís lifetime.

MI experts resumed their slipstacking study at 1:45 PM. They completed their study at 3:23 PM.

Operations terminated store 4347 at 6:15 PM. Operators made the CDF and D0 collision halls ready for access. Both experiments were out of access by 7:54 PM.

Operations put the TeV through a dry squeeze at 8:44 PM.

Operations put the TeV into shot setup at 9:16 PM.

Operations established combination store 4349 at 11:53 PM with an initial luminosity of 128.1E30.

At 11:55 PM, Operations reported that there was a problem with the Pbar sequencer that stopped Pbar from stacking. An operator called an expert.

Thursday, August 25
The midnight shift began with Operations monitoring store 4349, with Pbar resuming stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

Operations established beam to MCenter at midnight.

At 12:59 AM, a NuMI power supply (E:HV101) tripped off for the first of five times of the next six hours. The supply would eventually reset after each trip. After the last trip, Operations had an expert investigate.

Operations began preparing the Recycler for stashing antiprotons at 2:41 AM. The stash ended and Pbar resumed stacking at 3:21 AM. MCenter resumed taking beam five minutes later.

At 7:13 AM, MCenter beam stopped making it out of MI. An operator called a MI expert.

The Recycler began stashing at 7:50 AM. The stash ended and Pbar resumed stacking at 8:33 AM.

At 10:19 AM, MIRF5 and 18 tripped off while Booster experts were tuning their machine. Over the next twenty minutes MIRF5, 16, & 18 tripped off six times due to this tuning. The stations reset okay. An RF expert investigated.

At 12:06 PM, a Linac RF station (KRF1) tripped off for the first four times in forty minutes. It reset each time.

Electron Cooling experts opened up their enclosure MI-31 at 12:57 PM to make repairs on the Pelletron.

The Recycler began stashing at 12:58 PM.

At 1:52 PM, a NuMI power supply (HV101) tripped off. Operations contacted power supply experts.

The Recycler stash ended and Pbar resumed stacking at 2:08 PM.

Power supply experts had HV101 up and running by 2:19 PM.

MCenter resumed taking beam at 2:36 PM.

At 3:34 PM, a Booster quadrupole power supply (B:Q9) tripped off. An expert investigated and later had experts replace the supply.

At 4:10 PM, E:HV101 tripped off on a ground fault indication. Experts investigated and reported at 4:55 PM that the heat sink for the SCRs had a LCW leak. They had the supply up and running by 7:04 PM.

Friday, August 26
The midnight shift began with Operations monitoring store 4349 and the stash, with Pbar stacking, and with MiniBooNE and NuMI taking beam.

At 12:20 AM, KRF1 tripped off for the first of 10 times in a little over seven hours. It reset each time.

TeV expert began an end-of-store study at 7:07 AM.

The Plans for Friday and the Weekend
The plans for today are to complete the end-of-store study and then establish a new store. Pbar will hold back some antiprotons for a study after the store gets established. The rest of today and the weekend will consist of stacking, stashing, and storing.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/26/2005   email Fermilab